# CBVLM: Training-free Explainable Concept-based Large Vision Language Models for Medical Image Classification

### v1

@@ 5.5 Limitations
The main drawback of the proposed methodology for generating concept-based explanations with LVLMs lies in the independent prediction of each concept, which is a concern when applied to datasets with a large number of concepts. A similar problem arises with datasets with a large number of classes, as each class is provided as a separate option in the prompt. This can result in excessively long prompts that would require alternative ways of querying the model.
@@
Finally, CBVLM is only as good as the LVLM it uses. In a few cases, some of the LVLMs we tested struggled to provide answers to the questions (e.g., instead of answering the question, the LVLM simply described the input image), leading to an increase in unknown responses (as shown in Figures 3 and 6). This phenomenon is particularly common with SkinGPT, which may be due to its pretraining on very specific prompts designed to elicit desired responses [4]. LLava-Med also faced challenges in providing valid responses, especially in few-shot scenarios where $n>1$. This difficulty may stem from the fact that it is not trained with interleaved image-text data; in such models, image data is given at the input in the form of tokens, so as the number of shots increases, so does the number of tokens, making it more difficult for the model to pay attention to the full input sequence [52].
@@ 6 Conclusions and Future Work
In this supplementary material, we provide additional details that are not included in the main manuscript. Section A includes details of the HuggingFace checkpoints for each adopted Large Vision-Language Model (LVLM) in our experiments. Section B provides a list of the clinical concepts for each dataset. Section C presents the detailed prompts used in our experiments for both concept detection and disease diagnosis tasks, across all datasets.
Appendix A Implementation Details
@@
For SkinGPT-4 [4], we used the skingpt4_llama2_13bchat_base_pretrain_stage2 checkpoint available at https://github.com/JoshuaChou2018/SkinGPT-4.
@@
Idefics2 [54] HuggingFaceM4/idefics2-8b
@@
Table 3: HuggingFace checkpoints of the LVLMs used.
Appendix B Clinical Concepts
@@
Table 4: List of annotated clinical concepts of each dataset.
Appendix C Prompts
@@
Concept Detection.
The template prompts used to query LVLMs in the task of concept detection are given in Figures 9 to 11 for each dataset.
Disease Diagnosis.
The template prompts used to query LVLMs in the task of disease diagnosis are given in Figures 12 to 15 for each dataset.
@@
Figure 11: Example concept detection prompt for the SkinCon, CORDA and DDR datasets. This template can be applied to each of the 22 SkinCon concepts, 8 CORDA concepts, and 4 DDR concepts (see Table 4). The concept descriptions can be found in Figure 13. Only 1 demonstration is shown.
@@
Automatic Evaluation.
The template prompt used to query Mistral to perform the automatic extraction of the LVLMs’ responses is shown in Figure 16.
Sentence: <<<{sentence}>>>
Consider the sentence given in between <<<>>> and the following options: {options}
Choose the option that best fits the information conveyed by the sentence. Take your time and answer in json format by providing only the letter corresponding to the chosen option, following the template: {{‘Answer’: ‘option letter’}}. If the sentence does not provide enough information to choose an option, provide the following answer: {{‘Answer’: ‘UNK’}}.
Do not provide additional responses, context or explanations.
Answer:
Figure 16: Prompt to perform automatic extraction of the LVLMs’ responses with the Mistral LLM.

### v2

@@ 5.5 Limitations
The main drawback of the proposed methodology for generating concept-based explanations with LVLMs lies in the independent prediction of each concept, which is a concern when applied to datasets with a large number of concepts. A similar problem arises with datasets with a large number of classes, as each class is provided as a separate option in the prompt. This can result in excessively long prompts that would require alternative ways of querying the model. Although in medical applications the number of concepts is typically limited, a potential future direction to address concerns regarding the scalability for datasets with numerous concepts could be the adoption of tailored selection policies that prioritize only a small subset of the most informative concepts, as demonstrated in recent works [51].
@@
Finally, CBVLM is only as good as the LVLM it uses. In a few cases, some of the LVLMs we tested struggled to provide answers to the questions (e.g., instead of answering the question, the LVLM simply described the input image), leading to an increase in unknown responses (as shown in Figures 3 and 7). This phenomenon is particularly common with SkinGPT, which may be due to its pretraining on very specific prompts designed to elicit desired responses [4]. LLava-Med also faced challenges in providing valid responses, especially in few-shot scenarios where $n>1$. This difficulty may stem from the fact that it is not trained with interleaved image-text data; in such models, image data is given at the input in the form of tokens, so as the number of shots increases, so does the number of tokens, making it more difficult for the model to pay attention to the full input sequence [52].
Another relevant concern involves the robustness, security, and inherent biases of LVLMs [53]. While a comprehensive analysis of these aspects is beyond the scope of this work, it remains a critical consideration for future studies aimed at safe and reliable integration of LVLMs into clinical workflows. Collaboration between clinicians and AI researchers is crucial to ensure transparency and explainability of AI models deployed into clinical practice [54].
@@ 6 Conclusions and Future Work
In this supplementary material, we provide additional details that are not included in the main manuscript. Section A presents additional ablation studies comparing example selection strategies, namely RICES [24] and MMICES [25]. Section B offers further visualizations that illustrate the multidimensional interactions of concepts and model behavior in different LVLMs. Section C includes details of the HuggingFace checkpoints for each adopted Large Vision-Language Model (LVLM) in our experiments. Section D provides a list of the clinical concepts for each dataset. Section E presents the detailed prompts used in our experiments for both concept detection and disease diagnosis tasks, across all datasets.
Appendix A Comparison between example selection strategies - RICES and MMICES
In the main manuscript all experiments are conducted using Retrieval-based In-Context Examples Selection (RICES) [24] for the selection of the demonstration examples. This method selects the examples for in-context learning (ICL) using the similarity between visual features. Chen et al [25] extend RICES to also take into account text features, in a method called Mixed Modality In-Context Example Selection (MMICES), which the authors originally tested on vision-language datasets (e.g. image captioning, visual question answering). However, in our datasets such textual features do not exist. Thus, in the first stage of CBVLM (concept detection), MMICES cannot be applied. In the second stage (disease diagnosis), MMICES can be applied if we consider the concepts for each image as the textual features; the query image features are first compared to the train dataset image features, of which the $K$ ($K=200$) most similar are selected. Of these $K$ example images, we select the $n$ most similar to the query according to the cosine similarity between the concept vectors. The results of the comparison between the two approaches are presented in Figure 10, which shows the consistent improvement of RICES over MMICES. This might be explained by the fact that MMICES was proposed for tasks where the textual features have greater inter-image variability (e.g. different image captions for different images), while in our use-cases different images (especially from the same disease) might have the same set of concepts. Furthermore, we use a categorical representation of the set of concepts for each image, not a textual one. In the future it might be interesting to explore how to better adapt MMICES for this scenario or to use an image captioning model to provide textual descriptions of the images.
Refer to captionDerm7ptGenericMedical
Refer to captionSkinCon
Refer to captionCORDA
Refer to captionDDR
Figure 10: Disease diagnosis results with 4-shots and two example selection strategies (RICES [24] and MMICES [25]). In both methodologies, MedImageInsight [43] is the vision encoder used to extract the features from the query and demonstration images.
Appendix B Additional Visualizations
Figures 11 and 12 present alternative visualizations to Figures 3 and 7, respectively, in the form of spider plots, illustrating model behavior across different datasets for the tasks of concept detection and disease diagnosis. It is interesting to note that LVLMs trained on a given domain are able to, without training and with just a few demonstration examples, perform competitively in completely different domains (c.f. CheXagent on skin datasets or SkinGPT-4 on CORDA).
Figure 13 compares the per-concept performance of each evaluated LVLM in each of the four tested datasets, highlighting the LVLMs’ ability to detect individual clinical concepts. As mentioned in the main text, and again can be verified here for each LVLM, underrepresented concepts are the ones for which the F1-scores are lower.
Figure 11: Comparison of LVLM concept detection performance across datasets under different $n$-shot settings, reported in terms of Balanced Accuracy. Each $n$-shot setting is represented by a different color. The gray and pink dashed lines indicate the performance of CBM and CLAT, respectively.
Figure 12: Comparison of LVLM disease diagnosis performance across datasets under different $n$-shot settings, reported in terms of Balanced Accuracy. Each $n$-shot setting is represented by a different color. The gray and pink dashed lines indicate the performance of CBM and CLAT, respectively.
Figure 13: Comparison of per-concept F1-scores across datasets for each evaluated LVLM, illustrating the relative performance patterns of the models in detecting individual clinical concepts. Values in parentheses indicate the number of samples in the dataset for the corresponding concept.
Appendix C Implementation Details
@@
For SkinGPT-4 [4], we used the skingpt4_llama2_13bchat_base_pretrain_stage2 checkpoint available at: https://github.com/JoshuaChou2018/SkinGPT-4.
@@
Idefics3 [34] HuggingFaceM4/Idefics3-8B-Llama3
@@
Appendix D Clinical Concepts
@@
Table 4: List of annotated clinical concepts of each dataset.
@@
Appendix E Prompts
@@
Concept Detection
The template prompts used to query LVLMs in the task of concept detection are given in Figures 14 to 16 for each dataset.
Disease Diagnosis
The template prompts used to query LVLMs in the task of disease diagnosis are given in Figures 17 to 20 for each dataset.
@@
Figure 16: Example concept detection prompt for the SkinCon, CORDA and DDR datasets. This template can be applied to each of the 22 SkinCon concepts, 11 CORDA concepts, and 4 DDR concepts (see Table 4). Figures 1819, and 20, respectively. Only 1 demonstration is shown.
@@
Lung lesion refers to an abnormal area within the lung that may represent a nodule, mass, or area of infiltrate. It can be benign or malignant and may appear as a localized opacity or irregularity on imaging.
@@
Atelectasis is the partial or complete collapse of a part of the lung, leading to volume loss and increased opacity on radiographs, which can result from airway obstruction, compression, or surfactant deficiency.
@@
Pleural other refers to pleural abnormalities not classified as effusion or pneumothorax, including conditions such as pleural thickening, plaques, or calcifications.